# Reference Coverage Analysis of OpenAlex compared to Web of Science and Scopus

### 29 Jan 2024 v1

@@ Abstract
OpenAlex is a promising open source of scholarly metadata, and competitor to the established proprietary sources, the Web of Science and Scopus. As OpenAlex provides its data freely and openly, it permits researchers to perform bibliometric studies that can be reproduced in the community without licensing barriers. However, as OpenAlex is a rapidly evolving source and the data contained within is expanding and also quickly changing, the question naturally arises as to the trustworthiness of its data. In this empirical paper, we will study the reference and metadata coverage within each database and compare them with each other to help address this open question in bibliometrics. In our large-scale study, we demonstrate that, when restricted to a cleaned dataset of 16,788,282 recent publications shared by all three databases, OpenAlex has average reference numbers comparable to both Web of Science and Scopus. We also demonstrate that the comparison of other core metadata covered by OpenAlex shows mixed results, with OpenAlex capturing more ORCID identifiers, fewer abstracts and a similar number of Open Access information per article when compared to both Web of Science and Scopus.
@@ 1 Introduction
OpenAlex [1] was released on January 1st 2022 by OurResearch as a replacement for the discontinued Microsoft Academic Graph (MAG) and is offered as a fully open source of scholarly metadata, with all data, API information and code released to the public. As observed in the comparative study by Scheidsteger & Haunschild [2], not all aspects of the MAG were reproduced, as patents were not captured in OpenAlex. Aside from this exception, OpenAlex is effectively a continuation and expansion of the MAG.
@@
As a widely used open source repository of scholarly metadata, OpenAlex has previously been the subject of research, such as the review by Velez-Estevez et. al. [3], which comparatively analysed various APIs to bibliometric corpora, including API interoperability, characteristics and their use in research practice, and Akbaritabar et. al. [4] who released a working paper on the migration of scholars which included a comparative study between Scopus and OpenAlex, limited to the coverage of scholars in Western and non-Western countries.
At this early stage of its development, OpenAlex is a highly dynamic data source whose characteristics are not well studied.
In this research report, we present the results of some investigations into the metadata coverage of OpenAlex, following previous quantitative comparisons of citation coverage of the MAG such as Martín-Martín et al. [5], intersections of OpenAlex, WoS and Scopus, such as the study of the MAG against Scopus and other databases by Visser [6], journal coverage analysis of WoS, Scopus and Dimensions, such as the study of Singh et al. [7], and data completeness, such as Delgado-Quirós and Ortega’s smaller scale comparison of OpenAlex to other databases [8], to provide insights into the suitability of OpenAlex for bibliometrics in its current state.
@@ 1.1 Reference Coverage
As a first step we compare average reference counts between the three databases, whereby the basis of the comparison are the complete databases, subsets of publications with the document type ‘article’, and a shared sub-corpus of publications covered by all three databases. However, references can also be used for an assessment of the coverage, i.e. the proportion of relevant research publications that are included in the database and accessible to users for analysis [7]. An insufficient or biased coverage of the relevant literature should rule out the use of a database for a particular study.
@@
There are different ways to determine the coverage of such a database, for example, the comparison with external lists of relevant sources or publication lists of a sample of representative researchers of the studied fields. One relatively simple way to study literature coverage is calculating the internal reference coverage of a database as a whole or in relation to grouping characteristics, such as disciplines, the literature of particular countries or language communities.
@@
The internal coverage is the proportion of those cited references of a publication set which are themselves covered as source items in the database, out of all cited references in the set. We refer to these as source references and in contrast, to references to items that are not themselves indexed in the database as non-source references (or references to non-source items). As an example, suppose a publication set of interest cites 5000 references, including multiple citations of the same works, of which 2000 are covered by the same database, then the internal coverage of this literature would be 40%.
@@
A more comprehensive introduction to this concept and an analysis of the Science Citation Index’s (SCI) internal coverage of disciplines–as an estimator for disciplinary coverage–is available in [9, chapter 7] and [10]. The great advantage of this type of analysis is that one does not need any external data which may be difficult and costly to collect. This reliance on only the assessed data source itself is also the major disadvantage, as one is limited to the cited reference data as present in the assessed data source with all its contingencies. Therefore one cannot simply extrapolate from the coverage of cited literature to the coverage of literature segments that were never cited in the source data, possibly as a direct consequence of the source database’s selection criteria. These considerations show why internal reference coverage provides merely a partial and possibly source-biased measurement of coverage.
@@
Nevertheless, when comparing citation index databases, the differences in internal reference collection can be a useful guide when choosing the most suitable data source for a planned bibliometric study. There are no established guidelines for numerical values of coverage proportions required to allow reliable studies to be carried out. But, for example, Moed [9], analysed the combined ISI Citation Indexes (the predecessor of today’s Web of Science) and found that the coverage rate, which is the proportion of references from the 2002 source year that refer to ISI source journals, was highest for Molecular Biology and Biochemistry, at around 90%, followed by human-focused Biological Sciences, Chemistry, Clinical Medicine and Physics and Astronomy. It was vastly lower in the Arts and Humanities and intermediary in the Social Sciences, Mathematics and Engineering.
@@
When using this indicator to compare OpenAlex, WoS and Scopus, we are less interested in an evaluation in absolute values, but rather in assessing how OAL performs in comparison to the two established bibliometric databases. Internal reference coverage depends on the size and possibly the disciplinary profile of a database as well as the accuracy of its reference matching procedure. As OAL is actually much larger than Scopus and WoS (see Table 1) it could be expected that its internal reference coverage is at least not lower than in the latter databases.
@@ 1.2 Open Metadata
The increasing discussion surrounding the open availability of various types of scholarly metadata in bibliometrics is not limited to reference coverage, but expands to other metadata [11]. For instance, the Initiative for Open Abstracts (I4OA [^2] https://i4oa.org/) advocates open abstracts of scholarly works and calls on scholarly publishers to submit them to Crossref, a Digital Object Indentifier (DOI) registration agency [12]. However, coverage analyses of Crossref suggest that not all publishers provide open scholarly metadata to Crossref [13, 14]. Another example of essential metadata is the use of ORCIDs to persistently identify authors, helping bibliometricians not only to disambiguate author names, but also to interlink different data from different sources based on the ORCID. As open data sources are essential for OpenAlex, we will expand our analysis to compare abstracts and ORCID coverage at the journal level. Moreover, we will assess the coverage of open access status information between OpenAlex and the proprietary databases WoS and Scopus. In contrast to abstract and author information, all three databases use the same source, the open access discovery service Unpaywall, to retrieve open access status information [15].
@@ 2.1 Data
In this section, we describe the data used in this study, and the reasoning for our choices of restrictions and subsets of this data. To enable a fair comparison between OpenAlex, and WoS and Scopus, we have created a ‘Shared Corpus’ containing records common to all three datasets based on an exact DOI match, which have been published between 2015 and 2022, where the DOI is unique to the record in all three databases. In a further step, the references of the publications in the Shared Corpus are restricted to those published between 1996 to 2022.
@@
As the Scopus database mainly contains items from 1996 onwards (although since 2015, pre-1996 cited references and backfiles of major publishers have been added [^3] https://blog.scopus.com/posts/breaking-the-1996-barrier-scopus-adds-nearly-4-million-pre-1996-articles-and-more-than-83/), and WoS and OpenAlex have had no such restriction, to avoid bias in the computation of source reference counts and internal coverage we further restrict references to those items published between 1996 and 2022.
@@
We include in Table 1 a section on 2015-2022 to illustrate the influence of the time restriction to the size of the Shared Corpus, and to give context to the DOI matching and deduplication work described in Section 2.2.
@@
In the Scopus and WoS databases, pre-computed ‘total reference counts’, are delivered by the data providers Elsevier and Clarivate, whereas ‘source reference counts’ are computed for each record by our data provider FIZ Karlsruhe. Both databases contain all references of a given publication, regardless of whether they refer to items contained within or not contained within their databases, i.e. whether they are source and non-source references respectively, and without a time restriction. In contrast, in OpenAlex there only exist source references (see the OpenAlex documentation [1] [^4] https://docs.openalex.org/api-entities/works/work-object#referenced_works–apart from a smaller segment of references to deleted items) and a ‘source reference count’ has been calculated by FIZ Karlsruhe in our database. This fact explains the empty values for the average total reference counts in Table 3. We therefore have to relate this number to the source reference counts taken from WoS and Scopus.
@@
In Table 1 and Figure 1, we provide a summary of the records available in each data source and the intersection of the three data sources with each other, based on exact matching of DOIs, over the entire corpus, and restricted to records published between 2015 and 2022. Additionally, we provide information for the size of each corpus when restricted to records classified as ‘article’ to demonstrate that this does not substantially decrease the relative scale of OpenAlex to WoS and Scopus.
In the Appendix (Section 5.1) we provide additional information on the binary intersections of the three data sources and their proportion of the whole of each dataset in Table 5 and in Figure 3, a Venn diagram of the intersections of each database without the 2015-2022 publishing restriction.
@@
Figure 1: Venn diagram of the intersection of database records based on exact DOI match, published between 2015 and 2022
@@
It can be calculated from Table 1, that while the Shared Corpus, after DOI deduplication, contains 23.6% and 25.6% of the records in WoS and Scopus, and 6.9% of those in OpenAlex, it contains 41.1%, 35.8% and 31.7% of the references in the whole corpora of WoS, Scopus and OpenAlex respectively.
@@
The Shared Corpus, after DOI deduplication, contains 74.3% of the records in WoS published between 2015 and 2022, and 60.8% of the records in Scopus published between 2015 and 2022 and 21.8% of OpenAlex published between 2015 and 2022, based on an exact DOI match.
2.2 DOI Match Deduplication
@@
When constructing the Shared Corpus as described in Section 2.1, we excluded records without a DOI and records where more than one publication item is attributed to the same DOI - as we are virtually not able to decide which item is the correct one for a given DOI in the latter case. These removed duplicate records account for 39,481 additional removals from the 2015-2022 published restriction. This accounts for the difference between the size of the Shared Corpus and the intersection of the three databases between 2015 and 2022. We also briefly investigated the distribution of duplicate DOIs by record type in the whole corpora and the time segment 2015-2022, which we discuss in Section 4.
2.2.1 DOI Error Margins
@@
In Table 2 it can be seen that Scopus has a significantly larger number of DOIs with multiple records associated with it, and subsequently also a larger number of records with shared DOIs between 2015 and 2022.
As records with a shared DOI or those without a DOI were excluded from the DOI matching step in the construction of the Shared Corpus, we see that the greatest number of records excluded was from OpenAlex, then WoS and the least was from Scopus.
As records without a DOI are not matched in our analyses there may be a significant underestimation of the total overlap of the databases as portrayed in Figures 1 and 3, and Table 5. This error margin can be inferred from the data in Table 2.
@@ 2.3 Methodology
To evaluate the reference and source reference coverage of WoS and Scopus against OpenAlex, we first used the reported reference counts and pre-calculated source reference counts as described in Section 2.1. The average total reference count and source reference count was computed for: each database, for records marked as ‘article’ (or comprising the document type ‘article’ alongside other type markings, in the case of Scopus and WoS,) and for the Shared Corpus resulting from the DOI match (publication years 2015-2022). These numbers were replicated by counting the total number of references and records in each database and taking the ratio (‘references per record’), as reported in Table 4. Then for the final results, queries were created to calculate and average the number of references with reference publication year 1996 to 2022, and the number of references that are linked to source items and publication years 1996 to 2022.
To determine metadata coverage (as detailed in Section 3.3), we used the Shared Corpus as described in Section 2.1. Here, we focus on the level of journals. For abstracts, ORCID and Open Access (OA) status information, we aggregated the number of items in OpenAlex, WoS and Scopus by journal. For each journal, a record was counted if the desired metadata property was available. In the case of OA, we counted the item if the OA status was not marked as closed.
@@ 3 Results and Discussion
3.1 Source and Total Reference Coverage
The comparable performance of OpenAlex to the proprietary data sources is similarly reflected in Table 3, where a naïve averaging of the source reference count leaves OpenAlex looking comparatively poor at 7.6 references per record to the 16.9 or 18.7 of WoS and Scopus (and well behind the other databases’ average total reference count) when restricted to the 2015-2022 corpus shared by the three databases, OpenAlex proves competitive with a higher average source reference count than both WoS and Scopus. This could be explained by the fact that OpenAlex comprises of many–possibly less scientific–publications with comparatively short reference lists. The latter dynamic suggests that Scopus still has disadvantages due to its indexing start in 1996, which is evened out when all databases’ reference counts are limited to the time segment 1996-2022.
However, the slight advantage for OpenAlex is reversed when references are restricted to reference years 1996-2022, with Scopus outperforming OpenAlex, and WoS performing worst.

### 26 Feb 2024 v2

@@ Abstract
OpenAlex is a promising open source of scholarly metadata, and competitor to the established proprietary sources, the Web of Science and Scopus. As OpenAlex provides its data freely and openly, it permits researchers to perform bibliometric studies that can be reproduced in the community without licensing barriers. However, as OpenAlex is a rapidly evolving source and the data contained within is expanding and also quickly changing, the question naturally arises as to the trustworthiness of its data. In this empirical paper, we will study the reference and metadata coverage within each database and compare them with each other to help address this open question in bibliometrics. In our large-scale study, we demonstrate that, when restricted to a cleaned dataset of 16,788,282 recent publications shared by all three databases, OpenAlex has average source reference numbers and internal coverage, respectively, comparable to both Web of Science and Scopus. We also demonstrate that the comparison of other core metadata covered by OpenAlex shows mixed results, with OpenAlex capturing more ORCID identifiers, fewer abstracts and a similar number of Open Access information per article when compared to both Web of Science and Scopus.
keywords:
Bibliometrics, Open Scholarly Metadata, Citation Analysis, Open Abstracts, ORCID
@@ 1 Introduction
OpenAlex [14] was released on January 1st 2022 by OurResearch as a replacement for the discontinued Microsoft Academic Graph (MAG) and is offered as a fully open source of scholarly metadata, with all data, API information and code released to the public. As observed in the comparative study by [15], not all aspects of the MAG were reproduced, as patents were not captured in OpenAlex. Aside from this exception, OpenAlex is effectively a continuation and expansion of the MAG.
@@
As a widely used open source repository of scholarly metadata, OpenAlex has previously been the subject of research, such as the review by [17], which comparatively analysed various APIs to bibliometric corpora, including API interoperability, characteristics and their use in research practice, and [1] who released a working paper on the migration of scholars which included a comparative study between Scopus and OpenAlex, limited to the coverage of scholars in Western and non-Western countries. However, at this early stage of its development, OpenAlex is a highly dynamic data source whose characteristics are not well studied.
In this research report, we present the results of some investigations into the metadata coverage of OpenAlex, following previous quantitative comparisons of citation coverage of the MAG such as [10], intersections of OpenAlex, WoS and Scopus, such as the study of the MAG against Scopus and other databases by Visser [20], journal coverage analysis of WoS, Scopus and Dimensions, such as the study of [16], and data completeness such as Delgado-Quirós and Ortega’s smaller scale comparison of OpenAlex to other databases [2] and Färber’s tool for comparing author records between databases [4], to provide insights into the suitability of OpenAlex for bibliometrics in its current state.
@@ 1.1 Reference Coverage
References are of central importance for bibliometric databases, as matching them to their target items forms the basis for the calculation of citation metrics. As a first step, we compare average reference counts between the three databases, whereby the basis of the comparison are the complete databases, then subsets of publications with the document type ‘article’, and a shared sub-corpus of publications covered by all three databases. Furthermore, references can also be used for an assessment of the coverage, i.e. the proportion of relevant research publications that are included in the database and accessible to users for analysis [16]. An insufficient or biased coverage of the relevant literature should rule out the use of a database for a particular study.
@@
There are different ways to determine the coverage of a database, for example, the comparison with external lists of relevant sources or publication lists of a sample of representative researchers of the studied fields. One relatively simple way to study literature coverage is calculating the internal reference coverage of a database as a whole or in relation to grouping characteristics, such as disciplines, the literature of particular countries or language communities.
@@
The internal coverage is the proportion of those cited references of a publication set which are themselves covered as source items in the database, out of all cited references in the set. We refer to these as source references and in contrast, to references to items that are not themselves indexed in the database as non-source references (or references to non-source items). As an example, suppose a publication set of interest has 5000 references, including multiple references to the same works, of which 2000 are covered by the same database, then the internal coverage of this literature would be 40%.
@@
A more comprehensive introduction to this concept and an analysis of the internal coverage of disciplines–as an estimator for disciplinary coverage–is available in [11, Chapter 7] and [19]. The great advantage of this type of analysis is that one does not need any external data which may be difficult and costly to collect. This reliance on only the assessed data source itself is also the major disadvantage, as one is limited to the reference data as present in the assessed data source with all its contingencies. Therefore one cannot simply extrapolate from the coverage of cited literature to the coverage of literature segments that were never cited in the source data, possibly as a direct consequence of the source database’s selection criteria. These considerations show why internal reference coverage provides merely a partial and possibly source-biased measurement of coverage.
@@
Nevertheless, when comparing citation index databases, the differences in internal reference collection can be a useful guide when choosing the most suitable data source for a planned bibliometric study. There are no established guidelines for numerical values of coverage proportions required to allow reliable studies to be carried out. But, for example, [11] analysed the combined ISI Citation Indexes (the predecessor of today’s Web of Science) and found that the coverage rate, which is the proportion of references from the 2002 source year that refer to ISI source journals, was highest for Molecular Biology and Biochemistry, at around 90%, followed by human-focused Biological Sciences, Chemistry, Clinical Medicine and Physics and Astronomy. It was vastly lower in the Arts and Humanities and intermediary in the Social Sciences, Mathematics and Engineering.
@@
When using this indicator to compare OpenAlex, WoS and Scopus, we are less interested in an evaluation in absolute values, but rather in assessing how OpenAlex performs in comparison to the two established bibliometric databases. Internal reference coverage depends on the size and possibly the disciplinary profile of a database as well as the accuracy of its reference matching procedure. As OpenAlex is actually much larger than Scopus and WoS (see Table 1) it could be expected that its internal reference coverage is at least not lower than in the latter databases.
@@ 1.2 Open Metadata
The increasing discussion surrounding the open availability of various types of scholarly metadata in bibliometrics is not limited to reference coverage, but expands to other metadata [18]. For instance, the Initiative for Open Abstracts (I4OA [^2] https://i4oa.org/) advocates open abstracts of scholarly works and calls on scholarly publishers to submit them to Crossref, a Digital Object Identifier (DOI) registration agency. However, coverage analyses of Crossref suggest that not all publishers provide open scholarly metadata to Crossref [12, 8]. Another example of essential metadata is the use of ORCIDs to persistently identify authors, helping bibliometricians not only to disambiguate author names, but also to interlink different data from different sources based on the ORCID [6]. As open data sources are essential for OpenAlex, we will expand our analysis to compare abstracts and ORCID coverage at the journal level. Moreover, we will assess the coverage of open access status information between OpenAlex and the proprietary databases WoS and Scopus. In contrast to abstract and author information, all three databases use the same source, the open access discovery service Unpaywall, to retrieve open access status information [3].
@@ 2.1 Data
In this section, we describe the data used in this study, and the reasoning for our choices of restrictions and subsets of this data. To enable a fair comparison between OpenAlex, and WoS and Scopus, we have created a ‘Shared Corpus’ containing records common to all three datasets based on an exact DOI match, which have been published between 2015 and 2022, where the DOI is unique to the record in all three databases, i.e. there are no multiple records with the same DOI. In the course of database ingestion, it is ensured that publications only ever have one DOI. In a further step, the references of the publications in the Shared Corpus are restricted to those published between 1996 to 2022.
@@
As the Scopus database mainly contains items from 1996 onwards (although since 2015, pre-1996 cited references and backfiles of major publishers have been added [^3] https://blog.scopus.com/posts/breaking-the-1996-barrier-scopus-adds-nearly-4-million-pre-1996-articles-and-more-than-83), and WoS and OpenAlex have had no such restriction, to avoid bias in the computation of source reference counts and internal coverage we further restrict references to those items published between 1996 and 2022.
@@
We include in Table 1 a section on articles published 2015-2022 in all three databases, to illustrate the influence of the time restriction to the size of the Shared Corpus, and to give context to the DOI matching and deduplication work described in Section 2.2.
@@
In the Scopus and WoS databases, pre-computed total ’reference counts’, are delivered by the data providers Elsevier and Clarivate, whereas ’source reference counts’ are computed for each record by our data provider FIZ Karlsruhe. Both databases are expected to contain all references of a given publication, regardless of whether they refer to items contained within or not contained within their databases, i.e. whether they are source and non-source references respectively, and without a fixed time restriction. In contrast, in OpenAlex there only exist source references (see the OpenAlex documentation [14] [^4] https://docs.openalex.org/api-entities/works/work-object#referenced_works–apart from a smaller segment of references to supposedly deleted items) and a ‘source reference count’ has been calculated by FIZ Karlsruhe in our database. This fact explains the empty values for the average total reference counts in Table 3. We therefore have to relate this number to the source reference counts taken from WoS and Scopus.
@@
In Table 1 we provide a summary of the records available in each data source and in Figure 1 we provide a diagram of the intersections between the three data sources, based on exact matching of unique DOIs, over the entire corpus, and restricted to records published between 2015 and 2022. Additionally in Table 1, we provide information for the size of each corpus when restricted to records classified as ‘article’ to demonstrate that this does not substantially decrease the relative scale of OpenAlex to WoS and Scopus.
@@
Figure 1: Venn diagram of the intersection sizes of unique DOIs based in each database on exact DOI match, for records published between 2015 and 2022
@@
It can be calculated from Table 1, that while the Shared Corpus, after DOI deduplication, contains 23.6% and 25.6% of all records in WoS and Scopus, and 6.9% of those in OpenAlex, it contains 41.1%, 35.8% and 31.7% of the references in the whole corpora of WoS, Scopus and OpenAlex respectively.
@@
The Shared Corpus, after DOI deduplication, contains 74.3% of the records in WoS published between 2015 and 2022, and 60.8% of the records in Scopus published between 2015 and 2022 and 21.8% of OpenAlex published between 2015 and 2022.
2.2 DOI Match and Deduplication
@@
When constructing the Shared Corpus as described in Section 2.1, we excluded records without a DOI and records where more than one publication item is attributed to the same DOI – as we are virtually not able to decide which item is the correct one for a given DOI in the latter case. These duplicate records account for the removal of 39,481 publications (counted as distinct DOI) in addition to those resulting from the restriction to 2015-2022. This accounts for the difference between the size of the Shared Corpus and the nominal intersection of the three databases between 2015 and 2022.
2.2.1 Error Margins of the DOI Match
@@
Records with a shared DOI or those without a DOI were excluded from the DOI matching step in the construction of the Shared Corpus. In Table 2, which focuses on all publications in the three databases which are published between 2015 and 2022, it can be seen that Scopus has a significantly larger number of DOIs with multiple records associated with it. Altogether, OpenAlex has the greatest number of records without DOI, then WoS and Scopus.
As records without a DOI are not matched in our analyses there is a significant underestimation of the total size of the databases as portrayed in Figure 1, similarly records which have a shared DOIs are counted once. The error margin for the number of records compared to DOIs can be inferred from the data in Table 2.
Another reason for the exclusion of items in the DOI match, which at the same time restricts to publication years 2015-2022, is the fact that publication years are not always exactly the same between databases, possible due to differences in the handling of early access and print publication dates. We define the time restriction as applying to all three databases at the same time.
@@ 2.3 Methodology
To evaluate the reference and source reference coverage of WoS and Scopus against OpenAlex, we first used the reported reference counts and pre-calculated source reference counts as described in Section 2.1. The average total reference count and source reference count was computed for: each database, for records marked as ‘article’ (or comprising the document type ‘article’ alongside other type markings, in the case of Scopus and WoS) and for the Shared Corpus resulting from the DOI match (publication years 2015-2022). These numbers were replicated by counting the total number of references and records in each database and taking the ratio (‘references per record’), as reported in Table 4. Then for the final results, queries were created to calculate and average the number of references with reference publication year 1996 to 2022, and the number of references that are linked to source items and publication years 1996 to 2022.
To determine metadata coverage (as detailed in Section 3.3), we also used the Shared Corpus as described in Section 2.1. Here, we restrict to publication items published in journals. For this purpose, the publication type categorisations of Web of Science and Scopus were used and the OpenAlex publications were assigned to these via the DOI comparison of the Shared Corpus, so that OpenAlex could be compared bilaterally with the other two databases. We then specifically compared the coverage of abstracts, ORCIDs and Open Access (OA) status information by assessing whether items have (at least one) of these and aggregated by journal, that is, for each journal, a publication record was counted if the desired metadata property was available. In the case of OA, we counted the item if the OA status was not marked as closed. We have normalised the journal title to lowercase to aggregate the items.
@@ 3 Results and Discussion
3.1 Total and Source Reference Coverage
Table 3, in a naïve averaging of the source reference count, leaves OpenAlex looking comparatively poor at 7.6 references per record to the 16.9 or 18.7 of WoS and Scopus (and well behind the other databases’ average total reference count). However, when restricting to the 2015-2022 corpus shared by the three databases, OpenAlex proves competitive with a higher average source reference count than both WoS and Scopus. The fact that results vary greatly depending on the underlying corpus definition could be explained by the fact that OpenAlex comprises of many–possibly less scientific–publications with comparatively short reference lists. When focusing on the comparison of the average total reference counts between WoS and Scopus, it initially appears that Scopus outperforms WoS, however when considering records marked as articles they perform more comparably. This trend continues when observing the Shared Corpus and the Shared Corpus with references from 1996 to 2022. Notably here the difference between the source reference count and total reference count decreases as the restrictions are added. The results suggest that Scopus still has a small disadvantage due to its initial indexing start in 1996. Consequently, the slight advantage for OpenAlex is reversed when references are restricted to reference publication years 1996-2022, with Scopus outperforming OpenAlex, and WoS performing worst – however, differences are very small.